OpenStudy (anonymous):

Describe two ways you can determine if a point is a solution to a system of three equations.

OpenStudy (anonymous):

thanks all

OpenStudy (anonymous):

what ways do you know of?

OpenStudy (anonymous):

I don't really know..

OpenStudy (anonymous):

well how have you solved those in the past?

OpenStudy (anonymous):

By solving for y..

OpenStudy (anonymous):

they are certain terms...ill give you one, elimination..whats another that kinda goes along with that?

OpenStudy (anonymous):

I give

OpenStudy (anonymous):

@hartnn , i dont want to just give the answer so how can i help?

OpenStudy (anonymous):

in my opinion, Idk.

hartnn (hartnn):

by giving an example.

OpenStudy (anonymous):

okay, thanks so @NeedMathHelpNoa if you have a system of 3 equations such as x+y+z=3 x-y+2z=8 2x+3y-z=4 what are ways you would try to solve that.

hartnn (hartnn):

" if a point is a solution" assume a point

OpenStudy (anonymous):

that's news to me.. idk if I'd try to solve them? Like get the variables to other sides? or what..